HIGH-RESOLUTION MID-INFRARED SPECTROSCOPY OF VOLATILE ORGANIC COMPOUNDS USING LASER DIFFERENCE-FREQUENCY SPECTROMETER

\underline{W. CHEN}, E. POULLET, J. BURIE, D. BOUCHER, Laboratoire de Physicochimie d’Atmosphère, CNRS UMR 8101, Université du Littoral Cote d’Opale, 145 Av. Maurice Schumann, 59140 Dunkerque, France.

A continuous-wave (cw) high-resolution mid-infrared spectrometer has been developed based on difference-frequency generation (DFG) by mixing two cw autoscanned Ti :Sapphire lasers in GaSe crystal. The DFG spectrometer is continuously tunable in the wide spectral region of 8-20 μm without any ”mode hop”, with a linewidth of ∼ 1 MHz. DFG spectra of various volatile organic compounds, such as acetylene, ethylene, benzene, and toluene are investigated over the wavelength range from 10 to 15 μm with a resolution of ∼ 10^{−3} cm^{−1}. The present work is aimed at study of line parameters and its application to quantitative analysis of heavy molecules in the gas phase by vibrational absorption spectroscopy.